# Interaction imbalanced spin-orbit coupled quantum droplets

### v1

@@ IV.3 Spin-demixing and counterpropagating droplets after quenching the SOC wavenumber
Let us now turn our attention to the impact of the SOC wavenumber on the dynamical response of the two-component interaction imbalanced droplets both in the presence and absence of an external harmonic trap. Below, the droplet setting is initiated in each two-component ground state with parameters $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}=10$, $\delta g=0.1$, $k_{L}=0.5$, $\Omega=0.5$ and $N=50$. Focusing on the un-trapped case, we show the density evolution of each droplet component upon considering a quench to relatively small [Fig. 10(a), (b)] or large [Fig. 10(c), (d)] SOC wavenumbers quantified by $(k_{L})^{2}_{f}<\Omega$ and $(k_{L})^{2}_{f}>\Omega$ respectively. Independently of the final value of the SOC wavenumber, each droplet component experiences a momentum “kick" since $k_{L}$ is related to the system’s momentum, and hence the droplets become moving. In particular, in the case of $(k_{L})^{2}_{f}<\Omega$ depicted in Fig. 10(a), (b) we observe that both components feature a breathing oscillation while moving in the negative $x$-direction with a constant velocity. As such, we conclude that the droplet components remain mixed during the course of evolution. We also mention in passing that a relatively weak periodic population transfer among the components of maximum amplitude $\sim 6\%$ takes place (not shown) following the aforementioned droplets breathing.
@@ V Summary and Perspectives
We have investigated the ground state phases and the quench dynamics of interaction-imbalanced SOC droplets in one dimension. Our analysis is based on the suitable coupled system of extended Gross-Pitaevskii equations of motion in the presence and absence of an external harmonic trap. In the ground state, we find a variety of different droplet configurations bearing miscible characters among the components, and ranging from standard non-modulated to stripe patterns having either flat-top or Gaussian shapes. A key feature of our setup is the SOC, where, in particular, the Rabi-coupling favors population transfer between the involved spin states, which may thus experience atom number imbalance.
Specifically, the droplet components deform from flat-top to Gaussian distributions by increasing (decreasing) the intracomponent interaction ratio (total atom number). The spatially modulated droplet character depends on the SOC wavenumber. A higher SOC wavenumber (roughly larger than the Rabi coupling) promotes the formation of stripes atop the droplet. Spin imbalance increases with larger interaction strength, higher SOC wavenumber, or lower Rabi coupling. Conversely, increasing the Rabi coupling while keeping other parameters constant leads to a gradual transition from stripe to Gaussian, then to flat-top droplets, suppressing imbalance. In the stripe region, spin populations exhibit a nearly constant imbalance, which decreases in the Gaussian regime and becomes almost uniform in the flat-top regime.
@@
Turning to the dynamics we have first examined the behavior of the collective breathing motion of the two-component droplet system after a quench of the trap strength. This process induces a relatively small population transfer. It is found that the breathing mode frequency increases for larger intracomponent interaction ratio, while it features a maximum in terms of the SOC wavenumber at the location of the transition from a standard non-modulated to the stripe droplet. Additionally, a sudden reduction or increase of the Rabi-coupling is associated with minor intercomponent population exchange and the formation of fragmented moving droplets in the absence of the trap and oscillating ones with the trap frequency in its presence. On the other hand, following a quench of the SOC wavenumber leads to a drastically different dynamical response. In particular, independently of the postquench SOC wavenumber the droplets acquire finite momentum and feature a directed motion. In the absence of a trap, this motion refers to the same direction or opposite ones for small and large post-quench amplitudes respectively. This process is accompanied by spatial mixing or de-mixing of the droplet components which atop exhibit breathing and spatial delocalization respectively. The presence of the trap leads to similar characteristics but enforces an oscillatory droplet motion with the same frequency as the one of the trap. Droplets oscillate out-of-phase (in-phase) in the trap for larger (weak) quench amplitudes while being immiscible (miscible).
Based on our results, there is a multitude of intriguing extensions that may be followed in the future. An immediate step is to construct an effective potential picture in the case of large particle imbalance where the majority stripe droplet component plays the role of an external lattice potential for the minority one as was done in Ref. [37] but in the absence of SOC. Another interesting pathway is to calculate the excitation spectrum of these two-component droplet structures, especially focusing on the stripe configurations and their crossover to non-modulated droplets in the similar line as given in ref. [62, 63, 34]. It would worthwhile to pursue these droplet features in the ring geometry and detect the state of the droplet using the optomechanical couplings as developed in the recent works [64, 65]. Moreover, it would be worth examining the correlation patterns [59] associated with the stripe droplets, shedding light on the underlying two-body processes participating in the respective many-body states. Finally, the generalization of our setup to two dimensions in order to reveal the phases of two-component stripe droplets and their relation with supersolids [14] is desirable.
@@ Acknowledgements.
S.G. would like to acknowledge the financial support from the University Grants Commission - Council of Scientific and Industrial Research (UGC-CSIR), India. We also gratefully acknowledge our supercomputing facilities Param-Ishan and Param-Kamrupa (IITG), where all the numerical simulations were performed. The work of PM is supported by MoE RUSA 2.0 (Physical Sciences – Bharathidasan University). S. I. M is supported from the Missouri Science and Technology, Department of Physics, Startup fund. S. I. M acknowledges fruitful discussions with P. G Kevrekidis, G. C. Katsimiga and G. Bougas on the topic of droplets. R.R. acknowledges the postdoctoral fellowship supported by Zhejiang Normal University, China, under Grants No. YS304023964.

### v2

@@ IV.3 Spin-demixing and counterpropagating droplets after quenching the SOC wavenumber
Let us now turn our attention to the impact of the SOC wavenumber on the dynamical response of the two-component interaction imbalanced droplets both in the presence and absence of an external harmonic trap. Below, the droplet setting is initiated in each two-component ground state with parameters $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}=10$, $\delta g=0.1$, $k_{L}=0.5$, $\Omega=0.5$ and $N=50$. Focusing on the un-trapped case, we show the density evolution of each droplet component upon considering a quench to relatively small [Fig. 10(a), (b)] or large [Fig. 10(c), (d)] SOC wavenumbers quantified by $(k_{L})^{2}_{f}<\Omega$ and $(k_{L})^{2}_{f}>\Omega$ respectively. Independently of the final value of the SOC wavenumber, each droplet component experiences a momentum “kick" since $k_{L}$ is related to the system’s momentum [62], and hence the droplets become moving. In particular, in the case of $(k_{L})^{2}_{f}<\Omega$ depicted in Fig. 10(a), (b) we observe that both components feature a breathing oscillation while moving in the negative $x$-direction with a constant velocity. As such, we conclude that the droplet components remain mixed during the course of evolution. We also mention in passing that a relatively weak periodic population transfer among the components of maximum amplitude $\sim 6\%$ takes place (not shown) following the aforementioned droplets breathing.
@@ V Summary and Perspectives
We have investigated the ground state phases and the quench dynamics of interaction-imbalanced SOC droplets in one dimension. Our analysis is based on the suitable coupled system of extended Gross-Pitaevskii equations of motion in the presence and absence of an external harmonic trap. Particularly, a variety of different ground state droplet configurations are identified that host miscible components, and range from standard non-modulated to stripe patterns featuring flat-top or Gaussian density profiles. A key feature of our setup is the presence of SOC, which through the Rabi-coupling favors population transfer between the involved spin states and thus atom number imbalance.
It is found that droplets in both components deform from flat-top to Gaussian ones for increasing (decreasing) intracomponent interaction ratios (total atom number). On the other hand, the stripe (spatially modulated) droplet character is regulated by the SOC wavenumber. Spin imbalance increases with larger interaction strength, higher SOC wavenumber, or smaller Rabi coupling. Interestingly, an increase of the Rabi-coupling leads gradually to structural modifications from stripe to non-modulated Gaussian droplets and subsequently to flat-top ones whose imbalance is suppressed. This transition is also evident in the spin populations acquiring an almost constant imbalance in the stripe region and start decreasing in the Gaussian regime while tending to be almost the same when approaching the flat-top regime.
@@
Turning to the dynamics we have first examined the behavior of the collective breathing motion of the two-component droplet system after a quench of the trap strength. This process induces a relatively small population transfer. It is found that the breathing mode frequency increases for larger intracomponent interaction ratio, while it features a maximum in terms of the SOC wavenumber at the location of the transition from a standard non-modulated to the stripe droplet. Additionally, a sudden reduction or increase of the Rabi-coupling is associated with minor intercomponent population exchange and the formation of fragmented moving droplets in the absence of the trap and oscillating ones with the trap frequency in its presence. On the other hand, following a quench of the SOC wavenumber leads to a drastically different dynamical response. In particular, independently of the postquench SOC wavenumber the droplets acquire finite momentum and feature a directed motion. In the absence of a trap, this motion refers to droplet propagation in the same or opposite direction for small and large post-quench amplitudes respectively. This process is accompanied by spatial mixing or de-mixing of the droplet components which simultaneously exhibit breathing and spatial delocalization respectively. The presence of the trap enforces an oscillatory droplet motion characterized by the trap frequency. Out-of-phase (in-phase) droplet motion is evidenced for larger (weak) quench amplitudes wit the droplets being immiscible (miscible).
Based on our results, there is a multitude of intriguing extensions that may be followed in the future. An immediate step is to construct an effective potential picture in the case of large particle imbalance where the majority stripe droplet component plays the role of an external lattice potential for the minority one as was done in Ref. [38] but in the absence of SOC. Another interesting pathway is to calculate the excitation spectrum of these two-component droplet structures, especially focusing on the stripe configurations and their crossover to non-modulated droplets. [64, 65, 35]. Moreover, it would be worth examining the correlation patterns [47] associated with the stripe droplets, shedding light on the underlying two-body processes participating in the respective many-body states. An additional appealing direction would be to generalize the considered model in order to create linear droplet arrays similarly to the recent work of Ref. [66] in order to explore their transport or entanglement properties. The generalization of our setup in two dimensions in order to reveal the phases of two-component stripe droplets and their relation with supersolids [14] is desirable. Finally, it would be interesting to study the impact of optomechanical couplings [67, 68] in the identified droplet states.
@@ Acknowledgements.
S.G. would like to acknowledge the financial support from the University Grants Commission - Council of Scientific and Industrial Research (UGC-CSIR), India. We also gratefully acknowledge our supercomputing facilities Param-Ishan and Param-Kamrupa (IITG), where all the numerical simulations were performed. The work of PM is supported by MoE RUSA 2.0 (Physical Sciences – Bharathidasan University). S. I. M is supported from the Missouri University of Science and Technology, Department of Physics, Startup fund. S. I. M acknowledges fruitful discussions with P. G Kevrekidis, G. C. Katsimiga and G. Bougas on the topic of droplets. R.R. acknowledges the postdoctoral fellowship supported by Zhejiang Normal University, China, under Grants No. YS304023964.